OpenStudy (anonymous):

help me plzz Consider the line that passes through the point (-1, 6) and has a slope of -4. Part 1: Write the equation of this line using point-slope form. (2 points) Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. (2 points) Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. (2 points)

OpenStudy (anonymous):

@AccessDenied @arman1002 help plzz

OpenStudy (helder_edwin):

use the equation \[ \large y-y_0=m(x-x_0) \]

OpenStudy (helder_edwin):

u have the slope m=-4 and the point (-1,6)

OpenStudy (nurali):

Part 1 1) use point -slope form y-y1=m(x-x1) so the equation would look like y-6=-4(x+1) Part Two 2) then the equation becomes y-6=-4x-4 3) add six to both sides to get y=-4x+2 Part 3 4) use the slope and y-intercept of the above equation 5) the standard from would look like y-2=x+0

OpenStudy (nurali):

if correct please click best response.

OpenStudy (accessdenied):

Standard form is \(ax + by = c\). To get this form from \(y = \neg 4x + 2\), we simply add \(4x\) to each side. \(4x + y = 2\)